OpenStudy (jacky90):

HELP!! Will medal and fan The time it takes to cover the distance between two cities by car varies inversely with the speed of the car. The trip takes 8 hours for a car moving at 60 mph How long does the trip take for a car moving at 32 mph?

OpenStudy (sooobored):

let v- speed (velocity) let t - time let d- distance v*t= d or d/t = v

OpenStudy (skullpatrol):

The time it takes to cover the distance between two cities by car varies inversely with the speed of the car. $$\huge t= \dfrac{k}{v}$$ $$\huge v_1 • t_1 = v_2 • t_2$$ The trip takes 8 hours for a car moving at 60 mph. How long does the trip take for a car moving at 32 mph? $$\huge 60•8 = 32•t_2$$

OpenStudy (jacky90):

Thank you so much for the easy explanation SkullPatrol

OpenStudy (skullpatrol):

Thanks for asking :-)